# Standard Gaussian Process Can Be Excellent for High-Dimensional Bayesian Optimization

### v3

@@ B.1 GP Training
Second, we examined MAP training with alternative prior choices over the length-scale parameters. In the main paper, we use a diffuse prior, Uniform(0.001, 30). In practice, another commonly used choice is a concentrated Gamma prior, e.g., Gamma(3, 6), which is the default choice in the popular BO library BoTorch. To see how it is concentrated, in Fig. 10 we compare its probability density curve with other priors, including the aforementioned uniform prior and the diffuse Gamma priors over the amplitude and noise variance.
We found that employing a concentrated prior over the length-scale parameters will severely deteriorate the high-dimensional optimization performance. As shown in Fig 12 and 13, after placing the prior Gamma(3,6) over each length-scale parameter, both the ultimate and runtime optimization performance drops substantially across many tasks, such as Mopta08 (124), SVM (388) and Ackley(150, 150), no matter using ADAM or LBFGS to conduct training. To figure out the reason, we further investigated the surrogate learning performance. Following the same setting as in Fig. 3, we examined the prediction alignment after applying the prior Gamma(3,6) over the length-scale parameters. As depicted in Fig. 11 (third and fourth column), this concentrated prior leads to a severe prediction deviation from the ground-truth, indicating failure in learning the target function. This is consistent with the report in the SaasBO paper [6] (see Fig.1 of that paper). The quantitative results shown in Table 8 and 9 further confirm the substantially worse prediction accuracy as compared with GP training with a diffuse prior. In addition, we notice that the previous work[21] have also tested standard BO using Gamma(3,6) as the prior over the length-scale parameters. In their high-dimensional optimization experiments (see Fig. 5 of that paper), standard BO performs badly. We used our setting (SBO-1 and SBO-2) to re-run one experiment that maximizes a high-dimensional polynomial function. It turns out that both SBO-1 and SBO-2 can surpass the best method I-BNN reported by that paper.
Together these results have shown that a concentrated prior can severely limit GP’s capacity for high-dimensional function estimation, leading to inferior performance in high-dimensional optimization.
@@
Task MLE MAP-Uniform(1E-3, 30) MAP-Gamma(3, 6) MAP-Gamma(3, 6)
LBFGS ADAM
Mopta08(124) 0.034$\pm$0.006 0.035$\pm$0.005 0.837$\pm$0.112 0.856$\pm$0.114
Rover(60) 0.327$\pm$0.078 0.323$\pm$0.075 0.78$\pm$0.154 0.819$\pm$0.115
DNA(180) 0.149$\pm$0.038 0.166$\pm$0.049 1.014$\pm$0.097 1.026$\pm$0.101
SVM(388) 0.034$\pm$0.01 0.039$\pm$0.01 1.057$\pm$0.066 1.06$\pm$0.067
Ackley(150, 150) 0.017$\pm$0.008 0.026$\pm$0.007 0.019$\pm$0.004 0.019$\pm$0.005
Ackley(300, 150) 0.023$\pm$0.006 0.07$\pm$0.047 0.431$\pm$0.09 0.459$\pm$0.092
Hartmann6(300, 6) 0.017$\pm$0.013 0.178$\pm$0.325 0.881$\pm$0.045 0.906$\pm$0.034
Rosenbrock(100, 100) 0.084$\pm$0.013 0.081$\pm$0.012 0.9$\pm$0.156 0.925$\pm$0.172
Rosenbrock(300, 100) 0.09$\pm$0.064 0.116$\pm$0.078 0.971$\pm$0.222 1.001$\pm$0.217
Stybtang(200, 200) 0.043$\pm$0.018 0.054$\pm$0.024 0.698$\pm$0.05 0.732$\pm$0.034
Table 8: Mean Squared Error(MSE) averaged from five runs. The best results are bolded.
Task MLE MAP-Uniform(1E-3, 30) MAP-Gamma(3, 6) MAP-Gamma(3, 6)
LBFGS ADAM
Mopta08(124) 0.57$\pm$0.08 0.542$\pm$0.104 -1.337$\pm$0.058 -1.715$\pm$0.212
Rover(60) -0.975$\pm$0.216 -1.062$\pm$0.286 -$1$1.322$\pm$0.073
DNA(180) -0.256$\pm$0.261 -0.388$\pm$0.35 -1.426$\pm$0.054 -2.026$\pm$0.183
SVM(388) 0.581$\pm$0.119 0.504$\pm$0.128 -1.453$\pm$0.038 -2.095$\pm$0.129
Ackley(150, 150) 0.428$\pm$0.222 0.14$\pm$0.218 0.451$\pm$0.088 0.445$\pm$0.087
Ackley(300, 150) 0.232$\pm$0.155 -0.257$\pm$0.528 -1.006$\pm$0.091 -1.037$\pm$0.087
Hartmann6(300, 6) 1.2$\pm$0.169 0.67$\pm$0.867 -1.344$\pm$0.029 -1.618$\pm$0.128
Rosenbrock(100, 100) 0.64$\pm$0.081 0.606$\pm$0.1 -1.308$\pm$0.083 -1.598$\pm$0.42
Rosenbrock(300, 100) 0.518$\pm$0.225 0.347$\pm$0.325 -1.411$\pm$0.134 -1.967$\pm$0.413
Stybtang(200, 200) 0.574$\pm$0.164 0.421$\pm$0.228 -1.251$\pm$0.017 -1.41$\pm$0.067
Table 9: Test log likelihood averaged from five runs. The best results are bolded.
Figure 12: Ultimate optimization performance using a diffuse prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters. SBO-1 and SBO-2 use Uniform(0.001, 30).
Figure 13: Runtime optimization performance using a diffuse prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters. SBO-1 and SBO-2 use Uniform(0.001, 30).
B.2 Alternative Acquisition Functions
Next, we examined two other popular acquisition functions used in BO: Expected Improvement (EI) and Thompson sampling (TS). At each step, EI computes the expected improvement upon the best sample so far,
@@ B.2 Alternative Acquisition Functions
where $\Psi(\cdot)$ and $\Phi(\cdot)$ are the CDF and PDF of the standard Gaussian distribution, ${\bf x}^{*}$ is the best sample that gives the largest function value among all the queries up to the current step, $\mu({\bf x})$ and $\sqrt{v({\bf x})}$ are the posterior mean and standard deviation of $f({\bf x})$ given current training data. We used BOTorch to maximize EI through L-BFGS. To conduct TS at each step, we used Sobol Sequence [29] to sample 3K inputs in the domain, and jointly sampled the 3K function values from the GP posterior. We then selected the input with the largest function value as the next query. We denote this method by TS-3K. To explore more input candidates, we also employed the Lancos and conjugate gradient descent to approximate the posterior covariance matrix, and sampled 10K input candidates from Sobol Sequence. We denote this method as TS-10K. We used BOTorch implementation of both methods. We used ARD Matérn kernel and MLE training at each step.
The ultimate and runtime optimization performance for using EI, TS-3K and TS-10K, are shown in Fig. 14 and 15, respectively. As we can see, in most cases, such as Mopta08 (128), SVM (388), DNA (180) and Rosenbrock (100, 100), both EI and TS perform worse than UCB, and also worse than the best baseline. On Ackley(150, 150) and Ackley(300, 150), the performance of EI and TS are even among the worst.
@@
The limitation of EI might arise from the gradient vanishing during the optimization, as pointed by the recent work [1]. A remedy is to use log-EI as the acquisition function. The failure of the TS might be due to the high-dimension of the input space. A few thousand candidates might not be sufficient to cover the whole space; as a result, the selected queries are inferior in indicating the best direction to optimize the target function.
Figure 14: Ultimate performance with different acquisition functions. Note that SBO-1 and SBO-2 use UCB.
Figure 15: Runtime performance with different acquisition functions. Note that SBO-1 and SBO-2 use UCB.
B.3 Using Square Exponential (SE) kernels
Third, we considered another widely used kernel in BO, namely, the SE kernel. We tested with ARD SE and SE kernel using a single length-scale parameter. We used MLE training and UCB $(\lambda=1.5)$ to run BO. As depicted in Fig. 17 and 18, with smaller optimization dimensions, such as Mopta08 (124), NAS201 (3) and DNA (180), the performance of SE is among the best. However, with higher dimensions, e.g., above 200, the performance of SE drops substantially, such as in SVM (388) and Ackley(300, 150).
We argue that the inferior performance might be due to that the vanilla SE kernel is not powerful enough to enable GP to well capture high-dimensional functions. Consider a single length-scale parameter. we observe that the value of the SE kernel is determined by the squared distance between the inputs,
$\kappa_{\text{SE}}({\bf x},{\bf x}^{\prime})=a\exp\left(-\eta\|{\bf x}-{\bf x}^{\prime}\|^{2}\right).$ (7)
With the growth of the input dimension $d$, the kernel function value decays much faster than the increase of the distance: a squared decay coupling with the exponential decay. The learning can be easily driven to adapt to a sparse, diagonal-dominant covariance matrix, which roughly treats each data point as independent. This can make the learning fail to capture the correlation between function values, and result in inferior learning performance. By contrast, the value of the Matérn kernel (see (1)) is dominated by the distance rather than the squared distance. Therefore, with the growth of the distance, the kernel function value decays much slower than the SE kernel. It is easier for the optimization to adjust the length-scale parameter so as to avoid “diagonal-like” covariance matrices. Consequently, the correlation between function values can be better captured in the high dimensional input space.
To verify our argument, we examined the GP learning performance with ARD SE kernel and ARD Matérn kernel. We used Rosenbrock(d,d) and Hartmann6(d, 6) as test cases. The former is a “full-rank” function, namely, all the $d$ input variables are relevant to the function value. The latter has a “low-rank” structure in that only the first 6 variables out of the $d$ input variables determines the function value. For each test case, we increased the input dimension from 50 to 300. We randomly generated 500 training examples and 100 test examples. We then trained each method with MLE, and examined the mean square error (MSE) and the test log likelihood. We repeated the training and test procedure for five times, and report the average MSE, average test log likelihood, and the standard deviations in Fig. 16.
As we can see, in both test cases, when the input dimension $d$ does not exceed 150, the performance of using SE and Matérn-5/2 is close, in terms of both MSE and log likelihood. The two curves almost overlap. However, for $d>150$ in both test scenarios, the MSE significantly escalates when employing the SE kernel, accompanied by a substantial decrease in the test log likelihood. These observations suggest a failure of the SE kernel in effectively learning the target function. In comparison, employing the Matérn kernel maintains a consistently favorable prediction performance even when $d>150$. On the Rosenbrock dataset, the MSE gradually increases while the test log likelihood decreases, following a roughly linear trend. This outcome is expected as, with the rise in input dimension $d$, more training examples are necessary to sustain comparable performance. Furthermore, on the Hartmann6 dataset, little variation is observed in both MSE and test log likelihood with increasing the input dimension when using the Matérn kernel. This suggests that the GP has effectively discerned the pertinent input variables and mitigated the influence of irrelevant ones. In summary, these results collectively underscore the advantage of employing Matérn kernels for learning high-dimensional functions. It is worth noting that the most recent work [12] designs a log-normal prior over each length-scale parameter to overcome the learning issue with the ARD SE kernel.
Figure 16: Learning high-dimension functions with different kernels. We trained on 500 points, and tested on 100 points. Results shown with mean and standard deviation over 5 runs.
Figure 17: Ultimate optimization performance with different kernels. Note that SBO-1 and SBO-2 use Matérn-5/2 kernel.
Figure 18: Runtime optimization performance with different kernels. Note that SBO-1 and SBO-2 use Matérn-5/2 kernel.
B.4 Different Choice of $\lambda$ in UCB
Finally, we examined how the exploration parameter $\lambda$ in UCB (see (4)) influences the performance in high-dimensional optimization. To this end, we used ARD Maérn kernel and varied $\lambda$ from {0.5, 1.0, 1.5, 2.0}. We conducted MLE training via ADAM optimization. We show the ultimate and runtime optimization performance in Fig. 19 and 20, respectively. It can be seen that, in most benchmarks (Mopta08, Rover, NAS201, DNA, etc.), the performance with different $\lambda$’s is close. However, on DNA, SVM and Ackley, the performance of $\lambda=0.5$ is typically worse than the bigger choice of $\lambda$. This might be due to that those problems are more challenging (e.g., Ackley has many local minima), a larger exploration level ($\lambda$) is needed to identify better optimization outcomes.
Figure 19: Ultimate optimization performance with different exploration level $\lambda$ in UCB. Note that SBO-1 and SBO-2 use $\lambda=1.5$.
Figure 20: Runtime optimization performance with different exploration level $\lambda$ in UCB. Note that SBO-1 and SBO-2 use $\lambda=1.5$.

### v4

@@ C.2.2 MAP Training with Alternative Priors
Second, we examined MAP training with alternative prior choices over the length-scale parameters. In the main paper, we use a uniform prior, Uniform(0.001, 30). In practice, another commonly used choice is a concentrated Gamma prior, e.g., Gamma(3, 6), which is the default choice in the popular GP/BO libraries GPyTorch and BOTorch. To see how it is concentrated, in Fig. 11 we compare its probability density curve with other priors, including the aforementioned uniform prior and the diffuse Gamma priors over the amplitude and noise variance.
We observe that the prior distribution Gamma(3, 6) places a large probability mass around its mean 0.5, strongly regularizing the length-scale parameters toward this value. According to our theoretical analysis (see Section 4), this introduces a significant risk of gradient vanishing in high-dimensional settings, potentially leading to failure in GP training.
To verify the risk, we examined the alignment between the prediction and the ground-truth across the test examples from Mopta08 (124) and DNA (180). As depicted in Fig. 12 (third column), the concentrated prior leads to a severe prediction deviation from the ground-truth, indicating failure in learning the target function. This observation aligns with findings reported in the SaasBO paper [6] (see Fig.1 of that paper).
We further investigated the impact of this prior within the optimization loop. As shown in Fig. 9, after placing the prior Gamma(3,6) over each length-scale parameter, the BO performance drops substantially across many tasks, such as Mopta08 (124), SVM (388) and Ackley(150, 150), no matter using ADAM or LBFGS to conduct training.
@@
C.2.3 Fully Bayesian Training
Third, we investigated whether a fully Bayesian training approach would yield substantial improvements in BO. In particular, we can use MCMC to perform posterior inference of the GP parameters, and then integrate them out during the prediction. We conducted another round of BO experiments. For each step, we used NUTS to train GP with ARD Matérn. We adopted the same settings as used in SaasBO-NUTS [6]: where the mean parameter is assigned the standard Gaussian prior $\mathcal{N}(0,1)$, amplitude $\text{Gamma}(2,0.15)$, and noise variance $\text{Gamma}(0.9,10)$. We then assigned a uniform prior over each length-scale, $\text{Uniform}(0.001,30)$. The number of warm up steps was set to 512, and after the warm up stage, we ran 256 steps, and selected the samples at every 16-th step as the posterior samples. We show the running and final optimization performance in Fig 7 , and Fig. 8. We also show the average running time per optimization step in Table 4 and Table 5. It can be observed that while NUTS contributes to a slight improvement in several tasks, the enhancement is marginal.
However, running NUTS is much more computationally costly compared to doing MAP and MLE. For example, on DNA, NUTS takes 59x and 40x running time of MAP and MLE, respectively, on average in each step. Therefore, for practical purposes, using MAP or MLE for GP training appears sufficient. It is computationally efficient, and the sacrifice of optimization performance is minimal compared to much more costly training methods like NUTS.
C.3 Different Choice of $\lambda$ in UCB
Figure 13: Runtime optimization performance with different exploration level $\lambda$ in UCB.
We examined how the exploration parameter $\lambda$ in UCB influences the BO performance in high-dimensional optimization. To this end, we used the Maérn kernel and varied $\lambda$ from {0.5, 1.0, 1.5, 2.0}. We conducted MLE training via ADAM optimization. We show the ultimate and runtime optimization performance in Fig. 13, respectively. It can be seen that, in most benchmarks (Mopta08, Rover, NAS201, DNA, etc.), the performance with different $\lambda$’s is close. However, on DNA, SVM and Ackley, the performance of $\lambda=0.5$ is typically worse than the bigger choice of $\lambda$. This might be due to that those problems are more challenging (e.g., Ackley has many local minima), a larger exploration level ($\lambda$) is needed to identify better optimization outcomes.
C.4 Alternative Acquisition Functions
Figure 14: Final performance with different acquisition functions.
Figure 15: Runtime performance with different acquisition functions.
Finally, we evaluated three other acquisition functions for high-dimensional optimization: Expected Improvement (EI), log-EI and Thompson sampling (TS). At each step, EI computes the expected improvement upon the best sample so far,
@@ C.4 Alternative Acquisition Functions
where $\Psi(\cdot)$ and $\Phi(\cdot)$ are the CDF and PDF of the standard Gaussian distribution, ${\bf x}^{*}$ is the best sample that gives the largest function value among all the queries up to the current step, $\mu({\bf x})$ and $\sqrt{v({\bf x})}$ are the posterior mean and standard deviation of $f({\bf x})$ given current training data. We used BOTorch to maximize EI through L-BFGS. The most recent work [1] introduced log-EI to overcome the numerical challenges in optimizing the original EI acquisition function. Accordingly, we also tested log-EI using the BoTorch implementation. To conduct TS at each step, we used Sobol Sequence [28] to sample 3K inputs in the domain, and jointly sampled the 3K function values from the GP posterior. We then selected the input with the largest function value as the next query. We denote this method by TS-3K. To explore more input candidates, we also employed the Lancos and conjugate gradient descent to approximate the posterior covariance matrix, and sampled 10K input candidates from Sobol Sequence. We denote this method as TS-10K. We used BOTorch implementation of both methods. We used the Matérn kernel and MLE training at each step.
@@
The ultimate and runtime optimization performance for using EI, TS-3K and TS-10K, are shown in Fig. 14 and 15, respectively. As we can see, in most cases, such as Mopta08 (128), SVM (388), DNA (180) and Rosenbrock (100, 100), both EI and TS perform worse than UCB, and also worse than the best baseline. On Ackley(150, 150) and Ackley(300, 150), the performance of EI and TS are even among the worst. The failure of EI might arise from the gradient vanishing during the optimization, as pointed by the recent work [1]. The failure of the TS might be due to the high-dimension of the input space. A few thousand candidates might not be sufficient to cover the whole space; as a result, the selected queries are inferior in indicating the best direction to optimize the target function. The use of log-EI performed excellently, achieving results comparable to those obtained with UCB.
Appendix D Limitation
Our current work has identified gradient vanishing as a critical failure mode of SBO in high-dimensional settings. We have proposed an effective, robust initialization method to mitigate this issue. However, we have not yet analyzed whether there is an upper limit to the input dimension for which SBO can function properly, or what that limit might be. To answer these questions, we will continue our investigation in future work, both theoretically and empirically.